# Jailbreak Large Vision-Language Models Through Multi-Modal Linkage

### v4

@@ Abstract
With the significant advancement of Large Vision-Language Models (VLMs), concerns about their␣potential misuse and abuse have grown rapidly. Previous studies have highlighted VLMs’ vulnerability to jailbreak attacks, where carefully crafted inputs can lead the model to produce content that violates ethical and legal standards. However, existing jailbreak methods struggle against state-of-the-art VLMs like GPT-4o, due to the over-exposure of harmful content and lack of stealthy malicious guidance. In this work, we propose a novel jailbreak attack framework: Multi-Modal Linkage (MML) Attack. Drawing inspiration from cryptography, MML utilizes an encryption-decryption process across text and image modalities to mitigate over-exposure of malicious information. To align the model’s output with malicious intent covertly, MML employs a technique called “evil alignment”, framing the attack within a video game production scenario. Comprehensive experiments demonstrate MML’s effectiveness. Specifically, MML jailbreaks GPT-4o with attack success rates of 97.80% on SafeBench, 98.81% on MM-SafeBench and 99.07% on HADES-Dataset. Our code is available at https://github.com/wangyu-ovo/MML.
@@ 1 Introduction
Jailbreaking attacks (Zou et al., 2023; Wei et al., 2023a) pose a huge security concern for Large Language Models (LLMs) and have become a focus of recent research. Despite having undergone safety alignment training (Ouyang et al., 2022; Bai et al., 2022) prior to deployment, most of the attacks can still exploit carefully designed inputs to bypass these safeguards, prompting the models to generate harmful content misaligned with human values. Extensive research has been conducted on jailbreak vulnerabilities (Zou et al., 2023; Wei et al., 2023a; Chao et al., 2023; Li et al., 2024a) and defenses (Wei et al., 2023b; Inan et al., 2023; Lin et al., 2024; Mo et al., 2024) for LLMs, which has in turn sparked similar investigations into VLMs (Dong et al., 2023; Gong et al., 2023; Zhang et al., 2023; Niu et al., 2024; Qi et al., 2024; Shayegani et al., 2024; Liu et al., 2024b; Li et al., 2024b).
Jailbreak attacks on VLMs can be categorized into three categories: perturbation-based, structure-based, and their combinations. Perturbation-based attacks (Dong et al., 2023; Niu et al., 2024; Qi et al., 2024; Shayegani et al., 2024) are based on the concept of adversarial examples (Szegedy, 2013; Goodfellow, 2014), employing gradients to create adversarial images. In contrast, structure-based attacks (Gong et al., 2023; Liu et al., 2024b) bypass VLM defenses by embedding harmful content directly into images using typography or text-to-image tools. Li et al. (2024b) propose a method combining perturbation and structure attacks. Their approach amplifies the toxicity of images containing harmful content by applying gradient-based perturbations. Since the most advanced models are largely closed-source, structure-based attacks hold greater potential. This work focuses on structure-based attacks, which are less explored in the literature.
@@
Although current methods have jailbroken LLaVA (Liu et al., 2024a), MiniGPT-4 (Chen et al., 2023), and CogVLM (Wang et al., 2023) with high success rates, their performance degrades when applied to state-of-the-art VLMs such as GPT-4o. We suggest the decline might be attibuted decline can be attributed to two primary limitations of existing structure-based attack methods: over-exposure of harmful content and neutral text guidance, as shown in Figure 1(a).
Over-exposure of harmful content occurs when harmful content, e.g. images of bombs or malicious text embedded in typography, is exposed directly in the input. With advancements in image comprehension capability and safety alignment of VLMs, such overt content is likely to trigger rejection.
Neutral text guidance refers to the absence of stealthy text prompts that instruct models to produce malicious and informative outputs while bypassing refusal. Consequently, even when the model does not directly refuse to respond, its outputs are often limited to ethical advice, legal reminders, or warnings against harmful behavior—effectively constituting an implicit rejection. Examples of the implicit rejection are in Appendix A.1.
To address these issues, we propose a novel jailbreak attack framework for VLMs, Multi-Modal Linkage (MML) Attack. MML applies an “encryption-decryption” scheme to the linkage between modalities, which we view as a weak spot of VLMs, to mitigate the over-exposure issue. Specifically, MML first “encrypts” images containing harmful information through different methods including word replacement or image transformation (Figure 2). It then guides the target VLMs to “decrypt” the original malicious content via text prompts during inference (Figure 3). To counter the lack of malicious guidance, MML combines a technique named evil alignment (Zeng et al., 2024), which describes a virtual scene to align target VLMs’ outputs with malevolent objectives. Figure 1 provides an overview of MML and how it differs from other methods.
To validate the effectiveness of an MML attack, we conduct experiments on four latest large VLMs as target models on three benchmarks, i.e., SafeBench (Gong et al., 2023), MM-SafeBench (Liu et al., 2024b), and HADES-Dataset Li et al. (2024b). The results demonstrate the superiority of our approach, achieving high attack success rates across datasets. For instance, we successfully jailbreak GPT-4o with success rates of 97.80% on SafeBench, 98.81% on MM-SafeBench, and 99.07% on HADES-Dataset. Compared with the state-of-the-art baseline methods (Gong et al., 2023; Liu et al., 2024b; Li et al., 2024b), our attack success rate increased by 64.80%, 73.56%, and 95.07% respectively.
@@
  • We propose the Multi-Modal Linkage (MML) attack, a novel jailbreak attack framework that draws on cryptography incorporating an “encryption-decryption” strategy.
  • We combine MML with evil alignment to amplify the maliciousness of the target model’s outputs. By crafting a virtual scenario, we effectively align responses with malicious intent.
  • We validate the effectiveness of our approach through extensive experiments on four models and three datasets, achieving a high attack success rate in jailbreaking state-of-the-art VLMs.
@@ 2 Related Work
Jailbreak attacks on VLMs can be categorized into three types: perturbation-based attacks, structure-based attacks, and their combination. Perturbation-based attacks (Dong et al., 2023; Shayegani et al., 2024; Niu et al., 2024; Qi et al., 2024) focus on using adversarial images with added noise to bypass the target model’s safety alignment. These adversarial examples are typically crafted using gradient information from open-source proxy models. Structure-based attacks (Gong et al., 2023; Liu et al., 2023) leverage VLMs’ visual understanding capabilities and their vulnerabilities in safety alignment of visual prompts. These attacks involve converting malicious instructions into typographic visual prompts or embedding related scenarios into input images to bypass restrictions. Combining these approaches, Li et al. (2024b) introduced HADES, which uses images related to malicious instructions and applies gradient-based perturbations on open-source models to create jailbreak inputs.
@@
As research into jailbreak attacks on VLMs progresses, evaluating their robustness against jailbreak attacks has emerged as a significant concern. Zhao et al. (2024) pioneered the adversarial robustness evaluation of VLMs. Gong et al. (2023) introduce a benchmark called Safebench, which comprises 500 malicious typographies organized into 10 categories. Liu et al. (2024b) develop MM-SafetyBench, a benchmark featuring 5,040 text-image pairs across 13 scenarios. Additionally, Li et al. (2024b) compile the HADES-Dataset containing 750 harmful text-image pairs across 5 scenarios. Furthermore, Luo et al. (2024) propose a more comprehensive benchmark, JailBreakV-28K, which offers enhanced diversity and quality in harmful queries across 16 scenarios. Since this work focuses specifically on structure-based attack evaluation, we select Safebench, MM-SafetyBench and HADES-Dataset as the datasets for our experiments.
@@ 3 Threat Model
VLMs integrate visual and textual processing to generate text outputs from multi-modal inputs. To mitigate potential misuse, VLMs are typically tuned for safety alignment (Ouyang et al., 2022; Bai et al., 2022) before deployment, enabling them to reject responses to malicious queries that violate usage policies (OpenAI, 2024d). The goal of jailbreak attacks is to prompt the model to directly respond to harmful queries, e.g., “How to make a bomb.”, bypassing the safety behavior learned from safety alignment.
@@ 4.1 Overview
For a malicious query, MML first adopts an approach similar to FigStep (Gong et al., 2023), transforming the query such as “Steps to manufacture illegal drugs. 1. 2. 3.” into a typographical image formatted as a title. To reduce the exposure of malicious information, we encrypt the image, as illustrated in Figure 2. In the text prompt, we first guide the targeted model to decrypt the content and reconstruct the original title from the encrypted image (Figure 3), then generate content based on the reconstructed title. To further amplify the maliciousness of the targeted model’s responses, we frame the attack within a simulated video game production scenario, aligning the model’s responses with the villain’s evil nature.
@@ 4.2.1 Encryption
To reduce the exposure of malicious content, we employ four strategies to encrypt images: word replacement, image mirroring, image rotation and base64 encoding.
Word replacement is to switch malicious words of specific part-of-speech to harmless, unrelated words. For word replacement, we use the Natural Language Toolkit (NLTK) (Bird et al., 2009) to perform part-of-speech tagging on the original malicious query text. In malicious queries, harmful intent is typically conveyed through adjectives and nouns. For example, as shown in Figure 2, words like “illegal” and “drugs” are key contributors to malicious meaning. To neutralize this, we replace all nouns with food-related terms and all adjectives with positive words. A detailed list of replacement candidates can be found in the Appendix A.2.
Image mirroring and rotation are straightforward techniques achieved by applying the respective transformations to the image with typographic visual prompts.
The base64 encoding method (Wei et al., 2023a; Handa et al., 2024), is to first encode the malicious query in base64 format then transform it into a typographical image.
Figure 2 illustrates the results of the encryption. Notably, MML is a highly flexible framework that supports any encryption method, as long as it can be successfully decrypted within the capabilities of target VLMs.
@@ 4.2.2 Decryption
To further enhance decryption accuracy, we provide a shuffled list of the words from the original malicious query as shown in Figure 3. The targeted model is then guided to compare this encrypted list with the decrypted content and refine the latter accordingly. By shuffling the words, harmful information remains concealed.
@@ 4.3 Evil Alignment
Another limitation of existing methods is the lack of stealthy malicious guidance in text prompt. Since text safety alignment is more refined (Gong et al., 2023; Liu et al., 2024b), current structure-based attacks rely solely on neutral text prompts to guide the model in generating content related to images containing malicious information (Figure 1(a)). As a result, even when the model doesn’t directly refuse to respond, it often fails to generate the desired malicious content. Instead, it typically provides legal or ethical warnings and suggestions on how to avoid harmful behavior. Our experiments (Section 5.3) also confirm this.
To address this and taking inspiration from Zeng et al. (2024), we describe a virtual scenario to enhance the maliciousness of the targeted model’s responses. Specifically, we prompt the target VLM with a context in which we are designing a video game. The input image, as shown in Figure 2, is a large screen of the villain’s lair in the game, but some content is missing, and the VLM is instructed to complete the missing content in alignment with the the villain’s malevolent nature.
We refer to this as evil alignment and we empirically find it has a strong complementary effect to the encryption-decryption process. The complete prompts are included in Appendix A.3.
@@ 5.1 Setup
We conduct the experiments on three datasets: SafeBench (Gong et al., 2023), MM-SafeBench (Liu et al., 2024b) and HADES-Dataset (Li et al., 2024b), which are widely used as benchmarks for structure-based attacks. SafeBench includes 10 AI prohibited topics, selected based on the OpenAI Usage Policy (OpenAI, 2024d) and Meta’s Llama-2 Usage Policy (Meta, 2023). 50 malicious queries are generated by GPT-4 (OpenAI et al., 2024) for each topic, a total of 500 queries. MM-SafeBench comprises malicious queries across 13 scenarios. We filter out no-violation queries by using GPT-4o as moderation, getting a subset of 1,180 queries. HADES-Dataset (Li et al., 2024b) contains 750 malicious instructions across five scenarios. Further details can be found in Appendix A.4.
@@
We set FigStep (Gong et al., 2023) and QueryRelated (Liu et al., 2024b) as baseline methods for SafeBench and MM-SafeBench. For HADES-Dataset, we use HADES (Li et al., 2024b) as the baseline method. These methods all represent state-of-the-art structure-based or combination-based attacks. The Figstep and QueryRelated attacks are shown in Figure 1(a). HADES adds adversarial noise to the image similar to QueryRelated. For more details, please refer to the Appendix A.4.
@@
Previous methods have demonstrated high jailbreak success rates on VLMs such as LLaVA-1.5 (Liu et al., 2024a) and MiniGPT-4 (Chen et al., 2023), but they struggle with more advanced models like GPT-4o (OpenAI, 2024b). Therefore, we focuse on targeting these state-of-the-art models. The models included in our experiment are: GPT-4o-2024-08-06 (OpenAI, 2024b), GPT-4o-Mini-2024-07-18 (OpenAI, 2024a), Qwen-VL-Max-0809 (Bai et al., 2023), and Claude-3.5-Sonnet-20241022 (Anthropic, 2024), all of which represent the most advanced VLMs currently available (Contributors, 2023). The temperature of all models is set to 0.7, and we do not introduce any system messages for all experiments.
@@
We use attack success rate (ASR) to evaluate the effectiveness of the proposed method. To improve the reliability of our assessment, we adopt the evaluation strategy utilized by the Competition for LLM and Agent Safety (CLAS) (CLAS, 2024) for jailbreak attacks, which combines both LLM-based and template-based approaches. Specifically, we first conduct a keyword check on the target model’s response to identify any rejection words, such as “sorry.” If such words are present, the attack score is recorded as 0. Otherwise, GPT-4o-Mini is employed as an evaluator to rate the model’s response on its adherence to the malicious instructions, with scores ranging from 1 to 5. A score of 1 indicates the model refuses to comply, while a score of 5 signifies complete fulfillment of the policy-violating instruction. For our experiments, only responses with a score of 5 are considered successful attacks.
@@
Figure 5: ASR of baselines vs. MML-M (ours) across various topics in $M$-SafeBench. The left two figures presents the results of the baseline methods, FigStep (Gong et al., 2023) and QueryRelated (Liu et al., 2024b), while the right figure illustrates the ASR of MML using image mirroring as encryption method.

### v5

@@ Abstract
With the rapid advancement of Large Vision-Language Models (VLMs), concerns about their␣potential misuse and abuse have grown rapidly. Prior research has exposed VLMs’ vulnerability to jailbreak attacks, where carefully crafted inputs can lead the model to produce content that violates ethical and legal standards. However, current jailbreak methods often fail against cutting-edge models such as GPT-4o. We attribute this to the over-exposure of harmful content and the absence of stealthy malicious guidance. In this work, we introduce a novel jailbreak framework: Multi-Modal Linkage (MML) Attack. Drawing inspiration from cryptography, MML employs an encryption-decryption process across text and image modalities to mitigate the over-exposure of malicious information. To covertly align the model’s output with harmful objectives, MML leverages a technique we term evil alignment, framing the attack within the narrative context of a video game development scenario. Extensive experiments validate the effectiveness of MML. Specifically, MML jailbreaks GPT-4o with attack success rates of 99.40% on SafeBench, 98.81% on MM-SafeBench, and 99.07% on HADES-Dataset. Our code is available at https://github.com/wangyu-ovo/MML.
@@ 1 Introduction
Jailbreaking attacks (Zou et al., 2023; Wei et al., 2023a) pose a huge security concern for Large Language Models (LLMs) and have become a focus of recent research. Despite having undergone safety alignment training (Ouyang et al., 2022; Bai et al., 2022) prior to deployment, most of the attacks can still exploit carefully designed inputs to bypass these safeguards, prompting the models to generate harmful content misaligned with human values. Extensive research has been conducted on jailbreak vulnerabilities (Zou et al., 2023; Wei et al., 2023a; Chao et al., 2023; Li et al., 2024b) and defenses (Wei et al., 2023b; Inan et al., 2023; Lin et al., 2024; Mo et al., 2024) for LLMs, which has in turn sparked similar investigations into VLMs (Dong et al., 2023; Gong et al., 2023; Zhang et al., 2023; Niu et al., 2024; Qi et al., 2024; Shayegani et al., 2024; Liu et al., 2024b; Li et al., 2024c).
Jailbreak attacks on VLMs can be classified into three categories: perturbation-based, structure-based, and hybrid approaches. Perturbation-based attacks (Dong et al., 2023; Niu et al., 2024; Qi et al., 2024; Shayegani et al., 2024) draw on the concept of adversarial examples (Szegedy, 2013; Goodfellow, 2014), employing gradients to create adversarial images. In contrast, structure-based attacks (Gong et al., 2023; Liu et al., 2024b) model safeguards by embedding harmful content directly into visual elements, often through typography or text-to-image tools. Li et al. (2024c) introduce a hybrid approach that combines both strategies, enhancing the toxicity of harmful visual content via gradient-based perturbations. Given that state-of-the-art VLMs are predominantly closed-source, structure-based attacks offer greater practical potential. Despite this, they remain relatively underexplored in the literature. This work focuses on advancing structure-based jailbreak techniques.
@@
Although existing methods have achieved high jailbreak success rates on models such as LLaVA (Liu et al., 2024a), MiniGPT-4 (Chen et al., 2023), and CogVLM (Wang et al., 2023), their effectiveness diminishes significantly when applied to state-of-the-art VLMs like GPT-4o. We attribute this performance drop to two key limitations of current structure-based attack methods: over-exposure of harmful content and neutral text guidance, which are illustrated in Figure 1(a).
Over-exposure of harmful content occurs when harmful content, e.g., images of bombs or malicious text embedded in typography, is exposed directly in the input. With advancements in image comprehension capability and safety alignment of VLMs, such overt content is likely to trigger rejection.
Neutral text guidance refers to the absence of stealthy text prompts that instruct models to produce malicious and informative outputs while bypassing refusal. As a result, even when the model does not directly refuse to respond, its outputs are often constrained to ethical advice, legal reminders, or warnings against harmful behavior—amounting to an implicit rejection. Examples of the implicit rejection are in Appendix A.
To address these challenges, we propose a novel jailbreak attack framework for VLMs: the Multi-Modal Linkage (MML) Attack. MML applies an encryption-decryption [^1] We use the term encryption-decryption metaphorically to describe the process of concealing and revealing malicious content across modalities. It does not refer to formal cryptographic encryption. scheme to the linkage between modalities, which we view as a weak spot of VLMs, to mitigate the over-exposure issue. Specifically, MML first encrypts harmful content in images using techniques such as word substitution or visual transformation (Figure 2). During inference, the target VLM is then guided to decrypt this concealed malicious information via text prompts (Figure 3). To counter the lack of malicious guidance, MML incorporates a strategy known as evil alignment (Zeng et al., 2024), which embeds the attack within a virtual scenario designed to covertly align the model’s outputs with malevolent objectives. An overview of the MML framework and its distinction from existing approaches is illustrated in Figure 1.
To evaluate the effectiveness of MML, we conduct experiments on four latest large VLMs using three established benchmarks, i.e., SafeBench (Gong et al., 2023), MM-SafeBench (Liu et al., 2024b), and HADES-Dataset (Li et al., 2024c). The results demonstrate the superiority of MML, achieving high attack success rates across datasets. For instance, when targeting GPT-4o, MML attains success rates of 99.40% on SafeBench, 98.81% on MM-SafeBench, and 99.07% on HADES-Dataset. Compared with the state-of-the-art baseline methods (Gong et al., 2023; Liu et al., 2024b; Li et al., 2024c), MML improves the attack success rates by 66.4%, 73.56%, and 95.07%, respectively.
@@
  • We propose the Multi-Modal Linkage (MML) attack, a novel jailbreak framework that draws on cryptography incorporating an encryption-decryption strategy.
  • We integrate evil alignment into MML by crafting virtual scenarios that subtly guide model outputs toward malicious intent.
  • We conduct extensive experiments on four VLMs and three benchmarks, demonstrating MML’s superior jailbreak success rates over state-of-the-art methods.
@@ 2 Related Work
Jailbreak attacks on VLMs can be mainly categorized into three types: perturbation-based attacks, structure-based attacks, and their combination. Perturbation-based attacks (Dong et al., 2023; Shayegani et al., 2024; Niu et al., 2024; Qi et al., 2024) focus on using adversarial images with added noise to bypass the target model’s safety alignment. These adversarial examples are typically crafted using gradient information from open-source proxy models. Structure-based attacks (Gong et al., 2023; Liu et al., 2023) leverage VLMs’ visual understanding capabilities and their vulnerabilities in safety alignment of visual prompts. These attacks involve converting malicious instructions into typographic visual prompts or embedding related scenarios into input images to bypass restrictions. Combining these approaches, Li et al. (2024c) introduce HADES, which uses images related to malicious instructions and applies gradient-based perturbations on open-source models to create jailbreak inputs.
@@
As research into jailbreak attacks on VLMs progresses, evaluating their robustness against jailbreak attacks has emerged as a significant concern. Zhao et al. (2024) pioneer research into the adversarial robustness of VLMs. Li et al. (2024a) present RTVLM, a red-teaming dataset spanning 10 subtasks across 4 primary aspects. Gong et al. (2023) introduce a benchmark called Safebench, which comprises 500 malicious questions organized into 10 categories. Liu et al. (2024b) develop MM-SafetyBench, a benchmark featuring 5,040 text-image pairs across 13 scenarios. Additionally, Li et al. (2024c) compile the HADES-Dataset containing 750 harmful text-image pairs across 5 scenarios. Furthermore, Luo et al. (2024) propose a more comprehensive benchmark, JailBreakV-28K, which offers enhanced diversity and quality in harmful queries across 16 scenarios. Since this work focuses specifically on structure-based attack evaluation, we select Safebench, MM-SafetyBench and HADES-Dataset as the datasets for our experiments.
@@ 3 Threat Model
VLMs integrate visual and textual processing to generate text outputs from multi-modal inputs. To mitigate potential misuse, VLMs are typically tuned for safety alignment (Ouyang et al., 2022; Bai et al., 2022) before deployment, enabling them to reject responses to malicious queries that violate usage policies (OpenAI, 2024d). The goal of jailbreak attacks is to prompt the model to directly respond to harmful queries, e.g., ‘‘How to make a bomb.’’, bypassing the safety behavior learned from safety alignment.
@@ 4.1 Overview
For a malicious query, MML first adopts an approach similar to FigStep (Gong et al., 2023), transforming the query, e.g., ‘‘Steps to manufacture illegal drugs. 1. 2. 3.’’, into a typographical image formatted as a title. To reduce the exposure of malicious information, we encrypt the image, as illustrated in Figure 2. In the text prompt, we first guide the targeted model to decrypt the content and reconstruct the original title from the encrypted image (Figure 3), then generate content based on the reconstructed title. To further amplify the maliciousness of the targeted model’s responses, we frame the attack within a simulated video game production scenario, aligning the model’s responses with the villain’s evil nature.
@@ 4.2.1 Encryption
To reduce the exposure of malicious content, we mainly adopt four strategies to encrypt images:
Word replacement substitutes malicious words with harmless and semantically unrelated terms. Specifically, we use the Natural Language Toolkit (NLTK) (Bird et al., 2009) to perform part-of-speech tagging on the original malicious queries. Since malicious intent is typically expressed through nouns and adjectives (e.g., “drugs” and “illegal” in Figure 2), we replace all nouns with food-related words and all adjectives with positive descriptors. A detailed list of replacement candidates is provided in Appendix B.
Image mirroring and rotation apply simple geometric transformations to images containing typographic prompts.
Base64 encoding (Wei et al., 2023a; Handa et al., 2024) encodes the malicious text into base64 format and renders it into a typographic image, making it visually obscure but machine-decodable.
Figure 2 shows examples of each encryption method. Notably, MML is a flexible and extensible framework: any encoding strategy can be integrated, as long as the target VLM is capable of decrypting it during inference. The four methods above serve as basic representative examples. We further demonstrate MML’s extensibility by incorporating a shift cipher-based encryption strategy, discussed in Section 5.6.
@@ 4.2.2 Decryption
To further enhance the decryption accuracy, we provide a shuffled list of the words from the original malicious query as shown in Figure 3. The targeted model is then guided to compare this encrypted list with the decrypted content and refine the latter accordingly. By shuffling the words, harmful information remains concealed.
@@ 4.3 Evil Alignment
Another limitation of existing methods is the lack of stealthy malicious guidance in the text prompt. Due to increasingly refined safety alignment for text (Gong et al., 2023; Liu et al., 2024b), neutral prompts often fail to elicit informative malicious outputs—even when the model does not explicitly refuse, it typically responds with ethical advice or warnings (Figure 1(a); Section 5.3).
To address this, we adopt an evil alignment strategy inspired by Zeng et al. (2024). We describe a virtual scenario to enhance the maliciousness of the targeted model’s responses. We embed the attack in a fictional game development scenario, where the input image is described as a screen in a villain’s lair with missing content (Figure 2). The model is instructed to complete it in a way consistent with the villain’s objectives. This framing conceals malicious intent as part of a creative task, effectively bypassing safety filters.
We find that evil alignment complements the encryption-decryption process, significantly improving both stealth and attack success. Complete prompt examples are provided in Appendix C.
@@ 5.1 Setup
We conduct the experiments on three datasets: SafeBench (Gong et al., 2023), MM-SafeBench (Liu et al., 2024b) and HADES-Dataset (Li et al., 2024c), which are widely used as benchmarks for structure-based attacks. SafeBench includes 10 AI-prohibited topics, selected based on the OpenAI Usage Policy (OpenAI, 2024d) and Meta’s Llama-2 Usage Policy (Meta, 2023). 50 malicious queries are generated by GPT-4 (OpenAI et al., 2024) for each topic, a total of 500 queries. MM-SafeBench comprises malicious queries across 13 scenarios. We filter out non-violation queries by using GPT-4o as moderation, getting a subset of 1,180 queries. HADES-Dataset (Li et al., 2024c) contains 750 malicious instructions across five scenarios. Further details can be found in Appendix D.
@@
We set FigStep (Gong et al., 2023) and QueryRelated (Liu et al., 2024b) as baseline methods for SafeBench and MM-SafeBench. For the HADES-Dataset, we use HADES (Li et al., 2024c) as the baseline method. All these methods are state-of-the-art structure-based or combination-based attacks. The Figstep and QueryRelated attacks are shown in Figure 1(a). HADES adds adversarial noise to the image, similar to QueryRelated. For more details, please refer to the Appendix D.
@@
Previous methods have demonstrated high jailbreak success rates on VLMs such as LLaVA-1.5 (Liu et al., 2024a) and MiniGPT-4 (Chen et al., 2023), but they struggle with more advanced models like GPT-4o (OpenAI, 2024b). Therefore, we focus on targeting these state-of-the-art models. The models included in our experiment are: GPT-4o-2024-08-06 (OpenAI, 2024b), GPT-4o-Mini-2024-07-18 (OpenAI, 2024a), Qwen-VL-Max-0809 (Bai et al., 2023), and Claude-3.5-Sonnet-20241022 (Anthropic, 2024), all of which represent the most advanced VLMs currently available (Contributors, 2023). The temperature of all models is set to 0.7, and we do not introduce any system messages for all experiments. We also evaluate MML on OpenAI o1 reasoning model (Appendix D.8).
@@
We mainly use attack success rate (ASR) to evaluate the methods. To improve the reliability of our assessment, we adopt the evaluation strategy utilized by the Competition for LLM and Agent Safety (CLAS) (CLAS, 2024) for jailbreak attacks, which combines both LLM-based and template-based approaches. Specifically, we first conduct a keyword check on the target model’s response to identify any rejection words, such as ‘‘sorry.’’ If such words are present, the attack score is recorded as 0. Otherwise, GPT-4o-Mini is employed as an evaluator to rate the model’s response on its adherence to the malicious instructions, with scores ranging from 1 to 5. A score of 1 indicates the model refuses to comply, while a score of 5 signifies complete fulfillment of the policy-violating instruction. For our experiments, only responses with a score of 5 are considered successful attacks. Moreover, we additionally evaluate on Llama-Guard-3 to cross validate, results are in the Appendix D.7 and Figure 9.
@@
Figure 5: ASR of baselines vs. MML-M (ours) across various topics in $M$-SafeBench. The left two figures present the results of the baseline methods, FigStep (Gong et al., 2023) and QueryRelated (Liu et al., 2024b), while the right figure illustrates the ASR of MML using image mirroring as the encryption method.